OpenStudy (anonymous):

the manager of a coffee shop has one of coffee that sells for $6 per lb and another type that sells for $11 per lb.the manager wishes to mix 70lbs of the $11 coffee to get a mixture that will sell for $8 per lb. how many pounds of the $6 coffee should be used.

OpenStudy (amistre64):

amount@$6 = 70(11-8)/(8-6) i think ...

OpenStudy (anonymous):

i a getting good at these

OpenStudy (amistre64):

dbl chk 105 @ 6 :)

OpenStudy (anonymous):

put x = amount of $6 beans. then the total amount of beans will be 70 + x and you want that to cost $8 so the total cost will be \[8(70+x)\] that is one side of the equation. the other way to compute total cost is \[6x+70\times 11=6x+770\]

OpenStudy (anonymous):

so we have two expressions for total cost. set them equal and solve. should have \[6x+770=8(70+x)\] and i'll be damned if we don't get 105

OpenStudy (anonymous):

now i gotta figure out how sensei got it in one step, and i had to find an equation

OpenStudy (anonymous):

I had 87.5 at first but 105 was my second,thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok sensei what is your trick here?

OpenStudy (anonymous):

or must i figure it out for myself?

OpenStudy (anonymous):

i mean i see that solving my equation amounts to exactly the same thing. i am wondering if there is easy reasoning.

OpenStudy (amistre64):

my trick is to work it our in general and then add the specifics :) a+b = t; b = t-a ax + by = tz (t-b)x +by = tz tx -bx +by = tz b(y-x) = t(z-x) b@y = t(z-x)/(y-x) where; t=total; z=final; x=other; and y= ...well, y

OpenStudy (anonymous):

ok fine. my way is easier

OpenStudy (amistre64):

amount@$6 = 70(11-8)/(8-6) ... is another method; if I dont swap y and x I get the negative of it and can toss the sign

OpenStudy (anonymous):

i see what happened. you did all the algebra first that i would have to do second, so it is really the same.

OpenStudy (amistre64):

exactly :)

OpenStudy (amistre64):

so instead of hacking thru the same info time and again, I just use the general and fill in the blanks :)